# Distance Between Stochastic Linear Systems

### v1

@@ 4.4 Gap Metric Induced Type-$q$ Wasserstein Distance
$\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$ $\geq\mathbb{E}_{\pi_{\mathbf{f}}}\left[\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}-\sum^{2}_{i=1}\mathrm{Gap}_{i}(\theta_{i})\right)^{q}_{+}\right]$
@@
where we applied the Jensen’s inequality (for the convex function $f(x)=(x)^{q}_{+}$ with $q\geq 1$, $\mathbb{E}[(x)^{q}_{+}]\geq(\mathbb{E}[x])^{q}_{+}$) to get the second inequality. Taking infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ on both sides yields the desired result (59) and the proof is complete.
@@ 4.5 Numerical Demonstrations
We considered the following state space models as the nominal models for two different dynamical systems:
(60) $\Sigma_{1}$ $:A_{1}=\begin{bmatrix}0&1\\ -2&-0.5\end{bmatrix},B_{1}=\begin{bmatrix}0\\ 1\end{bmatrix},C_{1}=\begin{bmatrix}1&0\end{bmatrix},D_{1}=0$
(61) $\Sigma_{2}$ $:A_{2}=\begin{bmatrix}-3.2178&1.2354\\ -1.7812&-2.6507\end{bmatrix},B_{2}=\begin{bmatrix}0\\ 1\end{bmatrix},C_{2}=\begin{bmatrix}1&0\end{bmatrix},D_{2}=0.$
For generating the perturbed models, we choose $d=4$, $\theta_{1}\sim\mathcal{N}(0.01,0.01^{2})$, and $\theta_{2}\sim\mathcal{N}(0.05,0.05^{2})$. Totally, $N=50$ samples of perturbed models for each of the two dynamical systems were generated along the lines of (45). The proposed distance measure $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ between each of the models for both the systems was computed using (48) using linear programming based approach with the transport cost being the gap metric which was computed using the gapmetric command of Matlab. The upper bound and lower bounds for the proposed distance measure were computed using Proposition˜4.4 and Proposition˜4.8 respectively. The gap between the nominal models $\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}$ was found to be $0.7731$. We estimated the proposed distance measure $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}=0.7765$, and its lower and upper bounds as $0.6561$ and $0.8252$ respectively.
@@ 4.6 Summary
The proposed time domain distance given by (48) facilitates a simple linear programming based computation and also is devoid of additional supremum over frequency operation. We do not claim here that the upper and lower bounds given by Proposition˜4.4 and Proposition˜4.8 respectively are tight. In the next section, we will show that the frequency domain distance proposed in Section˜3 never exceeds the time domain distance proposed in this section.
@@ 5 Comparing Time Domain & Frequency Domain Distance Measures
It is well known from [25, 26] that $\nu$-gap can never exceed the gap metric for linear systems. Along those lines, we will prove in this section that the proposed frequency domain distance measure in Section˜3 between two stochastic LTI dynamical system never exceeds its time domain distance measure counterpart proposed in Section˜4. We formalise this observation using a comparison theorem followed by a simulation example based demonstration to corroborate our findings.
In the frequency domain setting, we know that the distribution $\mathbb{P}_{R_{i}(\omega)}$ governing the uncertainty of system $i$ with $i\in\{1,2\}$ on the boundary of the Riemann sphere is related to the distribution $\mathbf{f}_{\theta_{i}}$ of the random parameter $\theta_{i}$ through (19). Analogously, in the time domain setting, the distribution $\mathbf{f}_{i}$ of plant models of system $i\in\{1,2\}$ and the corresponding distribution $\mathbf{f}_{\theta_{i}}$ of the random parameter $\theta_{i}$ satisfy (43) and (44) with $\Phi_{i}$ denoting the measurable map from the parameter space to the state space of system plants as described earlier in Section˜4. We now define a transfer function mapping which when given a state space model, returns a real rational transfer function. That is, we define the transfer function mapping $\mathbf{TF}:\Phi_{i}\rightarrow\mathbf{R}\mathcal{L}_{\infty}$ such that $\mathbf{TF}(\Sigma_{i}(\theta_{i}))=P_{i}(\theta_{i};s)$. We also need a mapping $\Psi^{\mathfrak{R}}_{\omega}:\Phi_{i}\rightarrow\partial\mathfrak{R}$ that takes the state space model and maps it to the Riemann sphere after realising a real rational transfer function and subsequently evaluating it at a particular frequency $\omega$ and applying the inverse stereographic projection operation. Such a mapping can be defined using composition as
(62) $\Psi^{\mathfrak{R}}_{\omega}=\phi^{-1}\circ\mathcal{E}_{\omega}\circ\mathbf{TF}.$
Assumption 5.
@@
(63) $\pi_{\omega}:=\left(\Psi^{\mathfrak{R}}_{\omega}\times\Psi^{\mathfrak{R}}_{\omega}\right)_{\#}\pi_{\mathbf{f}}.$
Proof 5.2.
@@
$\pi_{\omega}(\mathbf{A}\times\partial\mathfrak{R})=\pi_{\mathbf{f}}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathbf{A})\times\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\partial\mathfrak{R})\right)$ $=\pi_{\mathbf{f}}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathbf{A})\times\Phi_{2}\right)$
@@
By similar arguments, we will also get $\pi_{\omega}(\partial\mathfrak{R}\times\mathbf{A})=\mathbb{P}_{R_{2}(\omega)}(\mathbf{A})$. Since we get the corresponding marginals, the result (63) follows immediately.
@@
(64) $\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)=\mathcal{R}_{P_{1}(\omega)}=\overline{\,\Psi^{\mathfrak{R}}_{\omega}\bigl{(}\mathrm{supp}\left(\mathbf{f}_{i}\right)\bigr{)}\,}.$
@@
(65) $\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\ =\ \Psi^{\mathfrak{R}}_{\omega}\bigl{(}\mathrm{supp}\left(\mathbf{f}_{i}\right)\bigr{)}.$
@@
(66) $\mathrm{supp}\left(\mathbf{f}_{i}\right)\subseteq\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}\left(\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\right).$
Proof 5.4.
To prove the forward inclusion of (64), let $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$ and let $r_{i}(\omega)=\Psi^{\mathfrak{R}}_{\omega}(\Sigma_{i})$. Take an open neighbourhood $\mathcal{B}_{r_{i}(\omega)}\subset\partial\mathfrak{R}$ such that $r_{i}(\omega)\in\mathcal{B}_{r_{i}(\omega)}$. By continuity of $\Psi^{\mathfrak{R}}_{\omega}$ via ˜5, there exists an open neighbourhood $\mathcal{B}_{\Sigma_{i}}$ of $\Sigma_{i}$ in the state space so that $\Psi^{\mathfrak{R}}_{\omega}(\mathcal{B}_{\Sigma_{i}})\subseteq\mathcal{B}_{r_{i}(\omega)}$. Since $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$, we have $\mathbf{f}_{i}(\mathcal{B}_{\Sigma_{i}})>0$, hence
@@
Therefore $r_{i}(\omega)\in\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$, proving $\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))\subseteq\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$. For the reverse inclusion, let $r_{i}(\omega)\in\partial\mathfrak{R}\setminus\overline{\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))}$. Then, there is an open neighbourhood $\mathcal{B}_{r_{i}(\omega)}$ of $r_{i}(\omega)$ such that $\mathcal{B}_{r_{i}(\omega)}\cap\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))=\emptyset$. By continuity of $\Psi^{\mathfrak{R}}_{\omega}$ through ˜5, we see that $\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})$ is open in the state space $\Phi_{i}$ and $\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\cap\mathrm{supp}\left(\mathbf{f}_{i}\right)=\emptyset$, hence $\mathbf{f}_{i}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\right)=0$. Thus $\mathbb{P}_{R_{i}(\omega)}(\mathcal{B}_{r_{i}(\omega)})=\mathbf{f}_{i}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\right)=0$, which shows $r_{i}(\omega)\notin\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$. So, $\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\subseteq\overline{\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))}$. Combining with the forward inclusion, the result (64) follows immediately. Additionally, if $\mathrm{supp}\left(\mathbf{f}_{i}\right)$ is compact, then $\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))$ is compact in $\partial\mathfrak{R}$ (due to continuity in ˜5), and hence it is closed too and thereby (64) reduces to (65). For proving (66), let $\mathbf{f}_{i}$ be a distribution on the state space with $(\Psi^{\mathfrak{R}}_{\omega})_{\#}\mathbf{f}_{i}=\mathbb{P}_{R_{i}(\omega)}$. If $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$, then by the above forward inclusion arguments, we have $\Psi^{\mathfrak{R}}_{\omega}(\Sigma_{i})\in\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$ which is equivalent to $\Sigma_{i}\in\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right))$. Hence, the result (66) follows.
@@
(67a) $\delta_{\nu}(P_{1},P_{2})$ $:=\sup_{\omega}\kappa(P_{1}(j\omega),P_{2}(j\omega)),\quad\text{where},$
(67b) $\kappa(P_{1}(j\omega),P_{2}(j\omega))$ $:=\mathrm{d_{chord}}(\phi^{-1}(P_{1}(j\omega)),\phi^{-1}(P_{2}(j\omega)))$
(67c) $=\frac{|P_{1}(j\omega)-P_{2}(j\omega)|}{\sqrt{(1+|P_{1}(j\omega)|^{2})(1+|P_{2}(j\omega)|^{2})}}$
denotes the pointwise-in-frequency gap between $P_{1}$ and $P_{2}$ and is exactly equal to the chordal distance between the two frequency-response points after inverse stereographic projection to $\partial\mathcal{R}$. Using this observation, we will now proceed ahead with the comparison theorem to formally establish the fact the proposed frequency domain distance that is described in Section˜3 through the definition given (3.14) never exceeds the proposed time domain distance counterpart that is described in Section˜4 through the definition given by (48).
@@
Consider two stochastic dynamical systems, whose state space models $\Sigma_{i}(\theta_{i})$ for system $i\in\{1,2\}$ are given by (40). Further, let the corresponding random transfer functions obtained from the respective state space models $\Sigma_{i}(\theta_{i})$ denoted by $P_{i}(\theta_{i};s)$ be given by (15). Additionally, let the random parameter $\theta_{i}$ affecting the system $i\in\{1,2\}$ be governed by the respective distribution $\mathbf{f}_{\theta_{i}}$. Subsequently, let the distribution $\mathbb{P}_{R_{i}(\omega)}$ governing the uncertainty of system $i$ with $i\in\{1,2\}$ on the boundary of the Riemann sphere be defined using (19) and the distribution $\mathbf{f}_{i}$ that governs the uncertainty of $\Sigma_{i}(\theta_{i})$ be given by (43). Then, for $q\geq 1$, we see that
(68) $\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}.$
Proof 5.6.
@@
(69) $\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})\ni\pi_{\mathbf{P}}:=(\mathbf{TF}\times\mathbf{TF})_{\#}\,\pi_{\mathbf{f}}.$
Further, for each frequency $\omega\geq 0$, the corresponding joint distribution on the Riemann sphere can be obtained using Lemma˜5.1 as
@@
$W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\big{)}$ $\leq\int_{\mathcal{R}_{P_{1}(\omega)}\times\mathcal{R}_{P_{2}(\omega)}}\mathrm{d_{chord}}(r_{1}(\omega),r_{2}(\omega))^{q}\,\pi_{\omega}(dr_{1},dr_{2})$
(70) $=\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\mathrm{d_{chord}}\big{(}\Phi^{\mathfrak{R}}_{\omega}(\Sigma_{1}),\Phi^{\mathfrak{R}}_{\omega}(\Sigma_{2})\big{)}^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2}).$
@@
(71) $\mathrm{d_{chord}}\big{(}\phi^{-1}(\mathcal{E}_{\omega}(P_{1})),\phi^{-1}(\mathcal{E}_{\omega}(P_{2}))\big{)}\ \leq\ \delta_{\nu}(P_{1},P_{2})\ \leq\ \delta_{g}(P_{1},P_{2}).$
Applying $P_{i}=\mathbf{TF}(\Sigma_{i})$ and the inequality (71) in (5.6), we get
$W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\big{)}$ $\leq\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}\!\big{(}\mathbf{TF}(\Sigma_{1}),\mathbf{TF}(\Sigma_{2})\big{)}^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
@@
$\underbrace{\sup_{\omega\geq 0}W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}}_{=\mathrm{d}_{q}(P_{1},P_{2})}\big{)}$ $\leq\sup_{\omega\geq 0}\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2})$

### v2

@@ 4.4 Gap Metric Induced Type-$q$ Wasserstein Distance
$\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
$\geq\mathbb{E}_{\pi_{\mathbf{f}}}\left[\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}-\sum^{2}_{i=1}\mathrm{Gap}_{i}(\theta_{i})\right)^{q}_{+}\right]$
@@
where we applied the Jensen’s inequality (for the convex function $f(x)=(x)^{q}_{+}$ with $q\geq 1$, $\mathbb{E}[(x)^{q}_{+}]\geq(\mathbb{E}[x])^{q}_{+}$) to get the second inequality. Taking infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ on both sides yields the desired result (60) and the proof is complete. ∎
@@ 4.5 Numerical Demonstrations
We considered the following state space models $(A_{1},B_{1},C_{1},D)$ and $(A_{2},B_{2},C_{2},D)$ as the nominal models for two different dynamical systems:
$A_{1}$ $=\begin{bmatrix}0&1\\ -2&-0.5\end{bmatrix},B_{1}=\begin{bmatrix}0\\ 1\end{bmatrix},C_{1}=\begin{bmatrix}1&0\end{bmatrix},D=0$ (61)
$A_{2}$ $=\begin{bmatrix}-3.2178&1.2354\\ -1.7812&-2.6507\end{bmatrix},B_{2}=\begin{bmatrix}0\\ 1\end{bmatrix},C_{2}=\begin{bmatrix}1&0\end{bmatrix}$. (62)
For generating the perturbed models, we choose $d=4$, $\theta_{1}\sim\mathcal{N}(0.01,0.01^{2})$, and $\theta_{2}\sim\mathcal{N}(0.05,0.05^{2})$. Totally, $N=50$ samples of perturbed models for each of the two dynamical systems were generated along the lines of (46). The proposed distance measure $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}$ between each of the models for both the systems was computed using (49) using linear programming based approach with the transport cost being the gap metric which was computed using the gapmetric command of Matlab. The upper bound and lower bounds for the proposed distance measure were computed using Proposition 4.2 and Proposition 4.4 respectively. The gap between the nominal models $\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}$ was found to be $0.7731$. We estimated the proposed distance measure $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}=0.7765$, and its lower and upper bounds as $0.6561$ and $0.8252$ respectively.
@@ 4.6 Summary
The proposed time domain distance given by (49) facilitates a simple linear programming based computation and also is devoid of additional supremum over frequency operation. We do not claim here that the upper and lower bounds given by Proposition 4.2 and Proposition 4.4 respectively are tight. In the next section, we will show that the frequency domain distance proposed in Section 3 never exceeds the time domain distance proposed in this section.
@@ 5 Comparing Time Domain & Frequency Domain Distance Measures
It is well known from [6, 7] that $\nu$-gap can never exceed the gap metric for linear systems. Along those lines, we will prove in this section that the proposed frequency domain distance measure in Section 3 between two stochastic LTI dynamical system never exceeds its time domain distance measure counterpart proposed in Section 4. We formalise this observation using a comparison theorem followed by a simulation example based demonstration to corroborate our findings.
In the frequency domain setting, we know that the distribution $\mathbb{P}_{R_{i}(\omega)}$ governing the uncertainty of system $i$ with $i\in\{1,2\}$ on the boundary of the Riemann sphere is related to the distribution $\mathbf{f}_{\theta_{i}}$ of the random parameter $\theta_{i}$ through (19). Analogously, in the time domain setting, the distribution $\mathbf{f}_{i}$ of plant models of system $i\in\{1,2\}$ and the corresponding distribution $\mathbf{f}_{\theta_{i}}$ of the random parameter $\theta_{i}$ satisfy (44) and (45) with $\Phi_{i}$ denoting the measurable map from the parameter space to the state space of system plants as described earlier in Section 4. We now define a transfer function mapping which when given a state space model, returns a real rational transfer function. That is, we define the transfer function mapping $\mathbf{TF}:\Phi_{i}\rightarrow\mathbf{R}\mathcal{L}_{\infty}$ such that $\mathbf{TF}(\Sigma_{i}(\theta_{i}))=P_{i}(\theta_{i};s)$. We also need a mapping $\Psi^{\mathfrak{R}}_{\omega}:\Phi_{i}\rightarrow\partial\mathfrak{R}$ that takes the state space model and maps it to the Riemann sphere after realising a real rational transfer function and subsequently evaluating it at a particular frequency $\omega$ and applying the inverse stereographic projection operation. Such a mapping can be defined using composition as
$\Psi^{\mathfrak{R}}_{\omega}=\phi^{-1}\circ\mathcal{E}_{\omega}\circ\mathbf{TF}.$ (63)
Assumption 5.1.
@@
$\pi_{\omega}:=\left(\Psi^{\mathfrak{R}}_{\omega}\times\Psi^{\mathfrak{R}}_{\omega}\right)_{\#}\pi_{\mathbf{f}}.$ (64)
Proof.
@@
$\pi_{\omega}(\mathbf{A}\times\partial\mathfrak{R})$ $=\pi_{\mathbf{f}}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathbf{A})\times\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\partial\mathfrak{R})\right)$
$=\pi_{\mathbf{f}}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathbf{A})\times\Phi_{2}\right)$
@@
By similar arguments, we will also get $\pi_{\omega}(\partial\mathfrak{R}\times\mathbf{A})=\mathbb{P}_{R_{2}(\omega)}(\mathbf{A})$. Since we get the corresponding marginals, the result (64) follows immediately. ∎
@@
$\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)=\mathcal{R}_{P_{1}(\omega)}=\overline{\,\Psi^{\mathfrak{R}}_{\omega}\bigl{(}\mathrm{supp}\left(\mathbf{f}_{i}\right)\bigr{)}\,}.$ (65)
@@
$\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\ =\ \Psi^{\mathfrak{R}}_{\omega}\bigl{(}\mathrm{supp}\left(\mathbf{f}_{i}\right)\bigr{)}.$ (66)
@@
$\mathrm{supp}\left(\mathbf{f}_{i}\right)\subseteq\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}\left(\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\right).$ (67)
Proof.
To prove the forward inclusion of (65), let $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$ and let $r_{i}(\omega)=\Psi^{\mathfrak{R}}_{\omega}(\Sigma_{i})$. Take an open neighbourhood $\mathcal{B}_{r_{i}(\omega)}\subset\partial\mathfrak{R}$ such that $r_{i}(\omega)\in\mathcal{B}_{r_{i}(\omega)}$. By continuity of $\Psi^{\mathfrak{R}}_{\omega}$ via Assumption 5.1, there exists an open neighbourhood $\mathcal{B}_{\Sigma_{i}}$ of $\Sigma_{i}$ in the state space so that $\Psi^{\mathfrak{R}}_{\omega}(\mathcal{B}_{\Sigma_{i}})\subseteq\mathcal{B}_{r_{i}(\omega)}$. Since $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$, we have $\mathbf{f}_{i}(\mathcal{B}_{\Sigma_{i}})>0$, hence
@@
Therefore $r_{i}(\omega)\in\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$, proving $\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))\subseteq\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$. For the reverse inclusion, let $r_{i}(\omega)\in\partial\mathfrak{R}\setminus\overline{\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))}$. Then, there is an open neighbourhood $\mathcal{B}_{r_{i}(\omega)}$ of $r_{i}(\omega)$ such that $\mathcal{B}_{r_{i}(\omega)}\cap\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))=\emptyset$. By continuity of $\Psi^{\mathfrak{R}}_{\omega}$ through Assumption 5.1, we see that $\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})$ is open in the state space $\Phi_{i}$ and $\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\cap\mathrm{supp}\left(\mathbf{f}_{i}\right)=\emptyset$, hence $\mathbf{f}_{i}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\right)=0$. Thus $\mathbb{P}_{R_{i}(\omega)}(\mathcal{B}_{r_{i}(\omega)})=\mathbf{f}_{i}\left(\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathcal{B}_{r_{i}(\omega)})\right)=0$, which shows $r_{i}(\omega)\notin\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$. So, $\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)\subseteq\overline{\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))}$. Combining with the forward inclusion, the result (65) follows immediately. Additionally, if $\mathrm{supp}\left(\mathbf{f}_{i}\right)$ is compact, then $\Psi^{\mathfrak{R}}_{\omega}(\mathrm{supp}\left(\mathbf{f}_{i}\right))$ is compact in $\partial\mathfrak{R}$ (due to continuity in Assumption 5.1), and hence it is closed too and thereby (65) reduces to (66). For proving (67), let $\mathbf{f}_{i}$ be a distribution on the state space with $(\Psi^{\mathfrak{R}}_{\omega})_{\#}\mathbf{f}_{i}=\mathbb{P}_{R_{i}(\omega)}$. If $\Sigma_{i}\in\mathrm{supp}\left(\mathbf{f}_{i}\right)$, then by the above forward inclusion arguments, we have $\Psi^{\mathfrak{R}}_{\omega}(\Sigma_{i})\in\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right)$ which is equivalent to $\Sigma_{i}\in\left(\Psi^{\mathfrak{R}}_{\omega}\right)^{-1}(\mathrm{supp}\left(\mathbb{P}_{R_{i}(\omega)}\right))$. Hence, the result (67) follows. ∎
@@
$\delta_{\nu}(P_{1},P_{2})$ $:=\sup_{\omega}\kappa(P_{1}(j\omega),P_{2}(j\omega)),\quad\text{where},$ (68a)
$\kappa(P_{1}(j\omega),P_{2}(j\omega))$ $:=\mathrm{d_{chord}}(\phi^{-1}(P_{1}(j\omega)),\phi^{-1}(P_{2}(j\omega)))$ (68b)
$=\frac{|P_{1}(j\omega)-P_{2}(j\omega)|}{\sqrt{(1+|P_{1}(j\omega)|^{2})(1+|P_{2}(j\omega)|^{2})}}$ (68c)
denotes the pointwise-in-frequency gap between $P_{1}$ and $P_{2}$ and is exactly equal to the chordal distance between the two frequency-response points after inverse stereographic projection to $\partial\mathcal{R}$. Using this observation, we will now proceed ahead with the comparison theorem to formally establish the fact the proposed frequency domain distance that is described in Section 3 through the definition given (31) never exceeds the proposed time domain distance counterpart that is described in Section 4 through the definition given by (49).
@@
Consider two stochastic dynamical systems, whose state space models $\Sigma_{i}(\theta_{i})$ for system $i\in\{1,2\}$ are given by (41). Further, let the corresponding random transfer functions obtained from the respective state space models $\Sigma_{i}(\theta_{i})$ denoted by $P_{i}(\theta_{i};s)$ be given by (15). Additionally, let the random parameter $\theta_{i}$ affecting the system $i\in\{1,2\}$ be governed by the respective distribution $\mathbf{f}_{\theta_{i}}$. Subsequently, let the distribution $\mathbb{P}_{R_{i}(\omega)}$ governing the uncertainty of system $i$ with $i\in\{1,2\}$ on the boundary of the Riemann sphere be defined using (19) and the distribution $\mathbf{f}_{i}$ that governs the uncertainty of $\Sigma_{i}(\theta_{i})$ be given by (44). Then, for $q\geq 1$,
$\mathrm{d}_{q}(P_{1},P_{2})\leq\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}.$ (69)
Proof.
@@
$\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})\ni\pi_{\mathbf{P}}:=(\mathbf{TF}\times\mathbf{TF})_{\#}\,\pi_{\mathbf{f}}.$ (70)
Further, for each frequency $\omega\geq 0$, the corresponding joint distribution on the Riemann sphere can be obtained using Lemma 5.1 as
@@
$W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\big{)}$
$\leq\int_{\mathcal{R}_{P_{1}(\omega)}\times\mathcal{R}_{P_{2}(\omega)}}\mathrm{d_{chord}}(r_{1}(\omega),r_{2}(\omega))^{q}\,\pi_{\omega}(dr_{1},dr_{2})$
$=\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\mathrm{d_{chord}}\big{(}\Phi^{\mathfrak{R}}_{\omega}(\Sigma_{1}),\Phi^{\mathfrak{R}}_{\omega}(\Sigma_{2})\big{)}^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2}).$ (71)
@@
$\footnotesize\mathrm{d_{chord}}\big{(}\phi^{-1}(\mathcal{E}_{\omega}(P_{1})),\phi^{-1}(\mathcal{E}_{\omega}(P_{2}))\big{)}\leq\delta_{\nu}(P_{1},P_{2})\leq\delta_{g}(P_{1},P_{2}).$ (72)
Applying $P_{i}=\mathbf{TF}(\Sigma_{i})$ & the inequality (72) in (5), we get
$W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}\big{)}$
$\leq\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}\!\big{(}\mathbf{TF}(\Sigma_{1}),\mathbf{TF}(\Sigma_{2})\big{)}^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
@@
$\underbrace{\sup_{\omega\geq 0}W_{q}^{q}\big{(}\mathbb{P}_{R_{1}(\omega)},\mathbb{P}_{R_{2}(\omega)}}_{=\mathrm{d}_{q}(P_{1},P_{2})}\big{)}$
$\leq\sup_{\omega\geq 0}\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2})$